# Detecting Masquerade Attacks in Controller Area Networks Using Graph Machine Learning

### v1

@@ Abstract
Modern vehicles rely on a myriad of electronic control units (ECUs) interconnected via controller area networks (CANs) for critical operations. Despite their ubiquitous use and reliability, CANs are susceptible to sophisticated cyberattacks, particularly masquerade attacks, which inject false data that mimic legitimate messages at the expected frequency. These attacks pose severe risks such as unintended acceleration, brake deactivation, and rogue steering. Traditional intrusion detection systems (IDS) often struggle to detect these subtle intrusions due to their seamless integration into normal traffic. This paper introduces a novel framework for detecting masquerade attacks in the CAN bus using graph machine learning (ML). We hypothesize that the integration of shallow graph embeddings with time series features derived from CAN frames enhances the detection of masquerade attacks. We show that by representing CAN bus frames as message sequence graphs (MSGs) and enriching each node with contextual statistical attributes from time series, we can enhance detection capabilities across various attack patterns compared to using only graph-based features. Our method ensures a comprehensive and dynamic analysis of CAN frame interactions, improving robustness and efficiency. Extensive experiments on the ROAD dataset validate the effectiveness of our approach, demonstrating statistically significant improvements in the detection rates of masquerade attacks compared to a baseline that uses only graph-based features, as confirmed by Mann-Whitney $\mathrm{U}$ and Kolmogorov-Smirnov tests $(p<0.05)$.
Index Terms:
Controller area networks, intrusion detection systems, graph ML, masquerade attacks.
@@ I Introduction
Controller area networks (CANs) have become ubiquitous in various industrial applications and in the automotive sector [1]. This protocol is essential for ensuring seamless communication among electronic control units (ECUs) that manage vital vehicle functions such as acceleration, braking, steering, and engine control [2, 3, 4]. The robustness, efficiency, and simplicity of CAN have established it as the industry standard for in-vehicle networks. However, the increasing connectivity and automation in modern vehicles, necessitating interfaces with external networks for diagnostics, firmware updates, and advanced driver assistance systems (ADAS), have exposed the inherent security vulnerabilities in CANs which lead to breach of integrity and confidentiality of in-vehicle communications [5, 6]. Among these, masquerade attacks stand out due to their stealth and potential for significant impact[7]. In such attacks, adversaries inject deceptive messages that mimic legitimate communication, manipulating vehicle behavior without triggering immediate detection [8]. The critical nature of these systems and the potential for severe consequences, such as unintended acceleration or braking, highlight the urgent need for robust IDS capable of identifying and effectively mitigating such sophisticated threats.
IDS have been widely employed to detect attacks such as fabrication, suspension, and the stealthiest of all, masquerade attacks, on CANs to ensure vehicle safety and functionality. Despite the effectiveness of IDS in detecting various types of attacks, CAN’s vulnerability to cyberattacks still poses significant risks [9]. Traditional IDS have primarily focused on signature-based or anomaly-based methods. Signature-based IDS rely on known attack patterns, which limits their effectiveness against novel threats [10, 11, 12, 13]. Anomaly-based IDs, while more flexible, often struggle with high false-positive rates and may lack the granularity needed to detect subtle, sophisticated attacks like masquerade attacks [14]. Masquerade attacks are particularly challenging because they involve injecting false messages that appear legitimate, allowing attackers to manipulate vehicle systems covertly [15]. Masquerade attacks usually bypass conventional detection mechanisms, making it crucial to develop more robust and precise IDS solutions [16]. Current methods often fall short in near real-time detection and handling of masquerade attacks, especially high-dimensional data typical of CANs [17, 18, 19, 20].
Detecting masquerade attacks in CANs is inherently challenging due to their ability to blend seamlessly into normal traffic. Traditional IDS approaches often struggle to identify these sophisticated intrusions because they do not cause immediate disruptions or obvious anomalies in the frequency of CAN frames [21]. Researchers have explored various methods to tackle this problem, focusing predominantly on time series analysis [22] and graph-based approaches [21]. Time series analysis leverages temporal patterns within CAN signals to detect anomalies, allowing for the identification of deviations in the sequence or timing of messages. However, while time series methods can pinpoint irregularities in CAN signal patterns, they often fall short in capturing the complex interactions between different ECUs and their signals. In contrast, graph-based methods have shown promise in addressing these complexities. By representing CAN bus data as graphs, these methods can model the intricate relationships and communication patterns between ECUs. This structural representation enables the detection of deviations from normal behavior that may indicate an attack, specifically in the case of fabrication attacks. Despite these advancements, current graph-based IDS approaches still face significant limitations. In particular, they often lack robustness in masquerade attack scenarios, failing to adequately detect these sophisticated intrusions due to their subtle nature. The primary limitations include insufficient incorporation of contextual data, such as temporal patterns and inter-signal relationships, which are crucial for accurate and reliable detection [21], [23]. Consequently, these limitations result in a higher likelihood of false negatives, where masquerade attacks go undetected, potentially leading to severe vehicle malfunctions or safety risks. Potential ways to address these limitations include integrating more comprehensive contextual data, exploring machine learning (ML) techniques that can better capture the complexities of CAN communications, and developing hybrid models that combine graph-based and time series-based approaches for enhanced detection accuracy.
The pressing need for an efficient approach that can accurately detect these stealthy attacks in resource-constrained vehicular environments forms the core motivation for this research. For instance, in a real-world scenario, a vehicle experiencing a masquerade attack might have its braking system manipulated without the driver’s knowledge, potentially leading to dangerous situations on the road. Detecting such an attack promptly could prevent accidents and enhance the overall safety of the vehicle. Our work aims to address these research gaps by leveraging graph ML to model CAN traffic using message sequence graphs [21] and time series analysis, enhancing the detection and analysis of masquerade attacks in the CAN bus.
@@
  • We propose a comprehensive framework for detecting masquerade attacks in the CAN bus. Our approach uses graph ML by integrating shallow graph embeddings and time series analysis to capture both the structural and temporal aspects of CAN frames. By representing CAN frames as message sequence graphs (MSGs), we enhance the ability to detect subtle masquerade attacks that traditional IDS might miss.
@@
  • We evaluate our proposed framework through extensive experiments using the ROAD dataset, a benchmark in vehicular network security research. Our results demonstrate significant improvements in detecting masquerade attacks when compared to an approach based only on graph topology. The evaluation highlights the scalability and robustness of our approach, ensuring its practical applicability in modern vehicular networks.
We have made the code available to reproduce all the results at [24].
@@ II Background
This section provides the necessary background on CAN and related topics. We begin with an overview of the CAN protocol (Section II-A) and its operational details (Section II-B). Next, we discuss the security aspects and common attack vectors in CAN (Section II-C). We then introduce the concept of graph-based representation in CAN, laying the foundation for how CAN communication patterns are modeled through MSGs (Section II-D), and explain the construction of MSGs (Section II-E). Finally, we cover the use of graph embeddings in the context of CAN (Section II-F).
II-A Controller Area Network (CAN)
CAN is a robust communication protocol that has become a standard in automotive and industrial applications for facilitating the exchange of messages between electronic control units (ECUs). Initially introduced by Robert Bosch GmbH, the protocol’s latest version (2.0) was released in 1991, marking a significant advancement in automotive communication technologies [25]. CAN was further standardized internationally in 2011 by ISO 11898-1:2011, which specified the requirements for high-speed communication and data exchange among various vehicle subsystems. CAN operates on a multi-master serial bus standard for ECUs, which allows devices to communicate with one another within a vehicle without a central computer. This protocol is designed to operate at high speeds, up to 1 Mbps, depending on the distance and configuration, with the ability to support distributed real-time communication efficiently [16].
II-B CAN Protocol
CAN operates as a broadcast protocol, enabling all ECUs connected to the bus to receive signals or messages transmitted over the bus. These signals, called frames, contain messages instructing the vehicle or system on executing operations. CAN encompasses the first two layers of the open systems interconnection (OSI) model, namely the physical and data-link layers, as shown in Fig. 1. Within CAN, ECUs such as the transmission control module, engine control unit, and the antilock braking system broadcast data frames that provide information about the vehicle’s current state.
@@ II-B CAN Protocol
Figure 1: CAN protocol frame structure with an 8-byte payload.
A CAN frame comprises several fields, each defined below along with their possible values:
  • Start of frame – This bit indicates the beginning of a CAN frame transmission — 1 bit, dominant (0). A dominant value ensures that all nodes on the network recognize the start of a new frame.
  • Identifier (ID) – This unique identifier represents the message priority, where lower values indicate higher priority — 11 bits for standard ID, 29 bits for extended ID. IDS are crucial for arbitration when multiple ECUs transmit simultaneously.
  • Remote transmission request (RTR) – Indicates whether the frame is for data transmission (dominant) or for requesting data (recessive) — 1 bit. Dominant (0) for data frames, recessive (1) for remote request frames. Dominant indicates an actual data frame, while recessive indicates a request.
  • Identifier extension bit (IDE) – Specifies whether the identifier is standard or extended — 1 bit. Dominant (0) for standard ID, recessive (1) for extended ID. This bit differentiates between the two ID formats.
  • Reserved – Reserved for future use and must be dominant to ensure compatibility — 1 bit. Must be dominant (0) but accepted as either dominant or recessive. Ensures consistency and compatibility across different CAN standards.
  • Data length code (DLC) – Indicates the length of the payload, specifying how many bytes of data are in the frame — 4 bits, representing 0–8 bytes. This is essential for the receiving ECU to know how much data to expect.
  • Payload – The actual data carried by the frame, encapsulated in bits — 0–8 bytes. Contains the information being transmitted, such as sensor readings or control commands.
  • Cyclic redundancy check (CRC) – Used to ensure the integrity of the transmitted data by detecting errors — 15 bits. Helps in error detection to ensure reliable communication.
  • Acknowledge (ACK) – Indicates whether the frame was acknowledged by any node on the network — 1 bit, recessive (1) when the frame is sent. A dominant value (0) signifies that at least one node successfully received the frame.
  • End of frame (EOF) – Denotes the end of a CAN frame — 7 bits, recessive (1). Marks the completion of the frame transmission.
  • Inter-frame spacing (IFS) – The space between the end of one frame and the start of the next — 3 bits, must be recessive (1). Ensures proper spacing between frames to avoid collisions.
In our research, we use only the ID and data fields. The ID serves as the message header that identifies the frame and is used for arbitration, the process of prioritizing frames when multiple ECUs transmit simultaneously, i.e., the lower the ID, the higher the priority. The data field holds the actual message contents of up to 8 bytes, where each unique information carried in the message is termed a signal. CAN frames with the same ID encode the same set of signals in the same format and are typically transmitted at a fixed frequency to relay updated signal values.
CAN protocol’s design ensures reliable communication in environments that are prone to electrical noise and disruptions, making it ideal for automotive applications where safety and efficiency are paramount [26]. Furthermore, the signal-level representation of CAN data is facilitated using a database for CAN (DBC) file, which translates binary payloads into real-valued signals, providing a time series representation of the network’s activity. This feature is crucial for the development of ML-based intrusion detection systems (IDS) for masquerade attacks as it captures a structured representation of CAN data, allowing to characterize the regular relationships between physical signals in a system [25, 27].
II-C CAN Security and Attacks
The security of CAN can be analyzed using the confidentiality, integrity, and availability (CIA) triad, which are fundamental principles of cybersecurity.
Confidentiality
CAN inherently lacks mechanisms for ensuring data confidentiality. Without encryption, all communications are susceptible to interception and interpretation. While some manufacturers have implemented cryptographic methods for specific functionalities like keyless entry [28], these are not standard features of the CAN protocol itself, leaving much of the network communication exposed.
Integrity
Although CAN employs a cyclic redundancy check (CRC) checksum to verify data integrity, this measure is insufficient for detecting malicious alterations. The CRC is designed to detect accidental alterations or corruptions, not intentional manipulations by authenticated sources such as the case of masquerade attacks. Thus, without an authentication mechanism, the protocol fails to guarantee the integrity of transmitted data [28].
Availability
The inherent design of CAN also impacts data and network availability. The arbitration rule allows higher-priority nodes to dominate communication and can be exploited to deny service to other nodes. While efficient under normal operations, this design choice becomes a liability when malicious actors manipulate the priority system to disrupt network access [28].
Our approach is designed to safeguard vehicles from various levels of cyberattacks, which can be broadly classified into three categories based on the attacker’s objectives [15]:
  • Fabrication attacks: A compromised ECU injects malicious IDs and data onto the CAN bus, manipulating the communication and potentially causing unintended behavior in the vehicle. All legitimate ECUs remain active and continue to send their original data. This type of attack is common and relatively easy to launch, as the attacker does not need to take control of any ECU [29].
  • Suspension attacks: An attacker disables a legitimate ECU, resulting in the disappearance of messages from the targeted ECU for a certain period. The attacker can achieve this by disconnecting the ECU from the in-vehicle network, thereby cutting off communication [30].
  • Masquerade attacks: These are the most sophisticated, stealthy, and destructive attacks in CAN. They combine elements of fabrication and suspension attacks, where the attacker silences a legitimate ECU and then impersonates it during ongoing operations to inject malicious messages at the expected frequency of benign messages [31].
In this research, we focus on masquerade attacks due to their intricate nature and their profound impact on in-vehicle networks. However, note that our framework will also be helpful for detecting fabrication and suspension attacks as they alter the regular patterns of frames and these changes are captured by the MSGs.
II-D Graph-Based Representation in CAN
A graph $G$ is defined as an ordered pair $G=(V,E)$, where $V$ represents a set of vertices (or nodes), and $E\subseteq(V\times V)$ represents a set of edges (or links) [32, 33]. Vertices act as entities and edges capture the relationships between nodes. For example, let us consider a toy graph $G$ with a vertex set denoted by $V(G)=\{v_{1},v_{2},v_{3},v_{4}\}$ and edge set by $E(G)=\{e_{1},e_{2},e_{3},e_{4},e_{5},e_{6}\}$. Fig. 2 depicts this graph for detailed visualization and to support the subsequent descriptions and definitions.
Figure 2: A representation of the nodes and edges of graph $G$.
II-D1 Directed graph
A directed graph, or digraph, consists of a vertex set $V$ and an edge set $E$ with two functions: $\textit{init}:E\to V$ and $\textit{ter}:E\to V$. These assign to each edge $e$ an initial vertex $\textit{init}(e)$ and a terminal vertex $\textit{ter}(e)$. The edge is considered directed from $\textit{init}(e)$ to $\textit{ter}(e)$. Directed graphs can include multiple edges between the same two vertices or loops, where an edge leads from a vertex to itself.
II-D2 Subgraph
A subgraph $G^{\prime}$ of a graph $G$ is defined by $V^{\prime}\subseteq V$ and $E^{\prime}\subseteq E$, denoted $G^{\prime}\subseteq G$. This implies $G$ includes $G^{\prime}$, or $G$ is a supergraph of $G^{\prime}$.
II-D3 Walks
A walk of length $k$ is an alternating sequence of vertices and edges in $G$, represented as $v_{0},e_{1},v_{1},e_{2},\ldots,e_{k},v_{k}$, where the edges $e_{i}={v_{i},v_{i+1}}$ for all $i<k$ [34, 35]. Here, $v_{0},v_{1},\ldots,v_{k}$ represent the vertices visited in order during the walk. A random walk is a particular case where each step from $v_{i}$ to $v_{i+1}$ is selected randomly, and which allows $v_{i}$ to be equal to $v_{i+1}$. The probability of a random walk that begins at node $i$ and ends up at node $j$ after $k$ steps can be represented by the transition matrix $p^{k}_{ij}$. Assuming that $G$ is a directed graph (digraph), and that the out-degree $\text{deg}^{+}(v_{j})$ of every vertex $v$ is greater than 0, we can define the transition matrix $P_{ij}$ as follows:
$P_{ij}:=\begin{cases}\frac{1}{\deg(v_{j})}&\text{if }(v_{i},v_{j})\text{ is anedge in the graph }G,\\ \frac{1}{\deg^{+}(v_{j})}&\text{if }(v_{i},v_{j})\text{ is an edge in the digraph }G,\\ 0&\text{otherwise.}\end{cases}$
II-E CAN Message Sequence Graph (MSG)
@@
The MSG is a crucial tool for analyzing the behavior of the CAN network [21]. It captures the sequence pattern of messages, reflecting their normal sequence of messages in CAN. Nodes in this graph correspond to unique CAN identifiers (IDs), denoting different ECUs or functions within the vehicle network. Edges map the sequence of messages between these IDS illustrating how information flows within the system. The MSG models regular patterns of communication between ECUs within specified sliding windows, allowing for detailed observation of both normal and potentially malicious activities. Typical sequence patterns found during various vehicle operations, such as parking or acceleration, can be captured by the MSG. These patterns form the basis for detecting deviations that may indicate cybersecurity threats like fabrication and suspension attacks.
Constructing an MSG involves parsing the CAN data to determine the sequence of CAN IDS and quantifying the frequency of these sequences. This process transforms the raw, time-stamped messages into a structured graph that highlights the dynamics of the network’s communication. For instance, typical sequences observed in vehicle operations—such as those related to increasing speed—involve a predictable order of CAN messages related to fuel delivery, RPM, and speed. Fig. 4 in Section IV provides a visualization of these sequences within the MSG. By focusing on the sequence and frequency of message exchanges, the MSG provides a powerful method for identifying disruptions in normal communication patterns indicative of potential attacks. This graph-based approach enhances the ability to scrutinize the integrated behavior of ECUs and ensure the integrity of vehicular communications.
II-F Graph Embeddings
Graph embeddings map nodes, edges, or entire graphs into a lower-dimensional vector space, preserving their structural and relational information [36]. Here, we focus on embedding the whole graph as our purpose is to detect intrusions in MSGs [37]. Consider $G=\{G_{1},\ldots,G_{m}\}$ as a collection of $m$ graphs, with each graph $G_{i}=\{V_{i},E_{i},l_{i}\}_{i=1}^{m}$ consisting of sets of vertices $V$ and edges $E$, and $l_{i}$ representing the class label of $G_{i}$. Whole-graph embedding aims to map an entire graph into a low-dimensional space $\mathbb{R}^{d}$ where $d\ll|V|$. This embedding vector should ideally retain as much of the graph’s structural information as possible [38].
In our analysis, we select node2vec[39], renowned for its effective network feature learning capabilities including node classification, link prediction, and community detection [39],[40]. Whole-graph embeddings are constructed by aggregating the node embeddings generated by node2vec, which captures the overall structure of the graph by summarizing the features of its individual nodes. This aggregation process typically involves averaging or pooling to combine node features into a single representation, providing a comprehensive view of the graph’s topology [41]. node2vec is particularly suitable for CAN bus data due to its proficiency in preserving the network topology and learning meaningful patterns of interactions between ECUs [42]. This is critical for identifying and classifying masquerade attacks within vehicular networks. The node2vec algorithm excels by extracting detailed feature representations for ECUs based on their communication patterns and roles within the network. node2vec parameters include walk length (l), which is crucial in determining the scope of each random walk, influencing whether the embeddings show immediate connectivity or broader network contexts. This adjustment allows feature learning to cover both local interactions and the entire network structure. The number of walks (r) initiated from each node deepens network exploration, ensuring a thorough representation by capturing diverse neighborhood configurations. Control over the random walks is further adjusted through the return parameter (p) and the in-out parameter (q). These parameters help direct the walk’s focus, switching between close neighborhoods and more distant connections. The return parameter (p) affects how likely the walk is to return to a node soon after departing, influencing the focus on local versus broad network connections. Conversely, the in-out parameter (q) shapes whether the walk tends to explore regions close to or far from the starting node, helping to balance the emphasis on tight community structures and a node’s role in the larger network topology. Additionally, the dimensions (d) parameter determines the size of the embedding vectors, balancing computational efficiency with the ability to capture rich graph-related properties.

### v2

@@ Abstract
Modern vehicles rely on a myriad of electronic control units (ECUs) interconnected via controller area networks (CANs) for critical operations. Despite their ubiquitous use and reliability, CANs are susceptible to sophisticated cyberattacks, particularly masquerade attacks, which inject false data that mimic legitimate messages at the expected frequency. These attacks pose severe risks such as unintended acceleration, brake deactivation, and rogue steering. Traditional intrusion detection systems (IDS) often struggle to detect these subtle intrusions due to their seamless integration into normal traffic. This paper introduces a novel framework for detecting masquerade attacks in the CAN bus using graph machine learning (ML). We hypothesize that the integration of shallow graph embeddings with time series features derived from CAN frames enhances the detection of masquerade attacks. We show that by representing CAN bus frames as message sequence graphs (MSGs) and enriching each node with contextual statistical attributes from time series, we can enhance detection capabilities across various attack patterns compared to using graph-based features only. Our method ensures a comprehensive and dynamic analysis of CAN frame interactions, improving robustness and efficiency. Extensive experiments on the ROAD dataset validate the effectiveness of our approach, demonstrating statistically significant improvements in the detection rates of masquerade attacks compared to a baseline that uses graph-based features only as confirmed by Mann-Whitney $\mathrm{U}$ and Kolmogorov-Smirnov tests $(p<0.05)$.
@@ I Introduction
Controller area networks (CANs) have become ubiquitous in modern vehicles and industrial applications, forming the backbone of communication between electronic control units (ECUs) responsible for essential functions such as acceleration, braking, and steering [halder2020coids, Jedh2021, shahriar_cantropy]. Renowned for their reliability and efficiency, CANs have become the de facto standard for in-vehicle communication. However, increased integration with external systems—such as diagnostics, firmware updates, and advanced driver assistance systems (ADAS)—has exposed CANs to a growing range of cyber threats that compromise data integrity and system functionality [Moriano:2022:Signal:Based:IDS]. CAN attacks are typically grouped into fabrication, suspension, and masquerade categories [verma2022addressing].
Among these threats, masquerade attacks are particularly insidious. These attacks inject malicious CAN frames that emulate legitimate messages, manipulating vehicle behavior while evading conventional intrusion detection systems (IDS) [Miller2015, sharmin2023comparative]. Unlike fabrication or suspension attacks, which create noticeable disruptions, masquerade attacks mirror the structure and frequency of authentic messages, making them exceptionally difficult to detect [verma2022addressing]. The consequences are severe: a compromised ECU could subtly trigger unintended acceleration, disengage braking mechanisms, or disable safety-critical features—actions that, if left unnoticed, can lead to catastrophic outcomes on the road.
While IDS solutions have been deployed to safeguard CANs, most fall into two categories: signature-based and anomaly-based approaches. Signature-based IDS are effective against known threats but fail against zero-day attacks [marfo2022condition, wang2024intrusion]. Anomaly-based IDS offer broader coverage but struggle with high rates of both false positives and negatives, as well as insufficient context-awareness for detecting stealthy behavior [sun2021anomaly, Xiao2023]. Existing solutions also tend to rely on either structural or signal-based analysis in isolation, failing to capture the critical relationship between message sequences and signal dynamics—a relationship that is essential for detecting masquerade attacks.
Masquerade attacks are uniquely challenging to detect due to their ability to seamlessly blend into normal traffic. These attacks mimic the frequency and content of legitimate frames, thereby bypassing traditional IDS techniques that rely on surface-level anomalies. Prior studies have shown that detection requires examining multiple consecutive frames, as single-message checks are often insufficient [Jedh2021]. Moreover, methods that use either graph structure or signal statistics alone often miss subtle patterns that span both temporal and structural domains [moriano2024benchmarkingunsupervisedonlineids, hanselmann2020canet]. For example, a vehicle under attack may gradually exhibit abnormal behavior—such as slowing to a halt—not due to one erroneous message, but due to a sequence of seemingly valid ones. Detecting such behavior demands a multimodal perspective.
To address these limitations, this work proposes a unified framework that integrates graph-based and time series analysis for the detection of masquerade attacks in CANs. By modeling CAN traffic as MSGs and enriching each node with contextual time series features, our method captures both structural irregularities and signal-level deviations. This dual perspective enables robust detection of subtle intrusions that would otherwise go unnoticed in traditional systems. We also report the testing time per window (TTW) [nichelini2023canova] as a proxy for inference time to assess the real-time detection capability of our approach.
@@
  • We propose a comprehensive framework for detecting masquerade attacks in the CAN bus. Our approach uses graph ML by integrating shallow graph embeddings and time series analysis to capture both the structural and temporal aspects of CAN frames. By representing CAN frames as MSGs, we enhance the ability to detect subtle masquerade attacks that traditional IDS might miss.
@@
  • We evaluate our proposed framework through extensive experiments using the ROAD dataset, a benchmark in vehicular network security research. Our results demonstrate significant improvements in detecting masquerade attacks when compared to an approach based graph topology only. The evaluation highlights the scalability and robustness of our approach, ensuring its practical applicability in modern vehicular networks.
We have made the code available to reproduce all the results at [GraphML-CONTROLLER-AREA-NETWORK-Attack-Detection].
@@ II Background
II-A CAN Protocol
CAN is a widely adopted message-based communication protocol designed to enable efficient data exchange among ECUs in vehicles and industrial systems [DiNataleZengGiustoGhosal2012CAN]. CAN operates on the physical and data-link layers of the OSI model, using a multi-master, broadcast-based communication model over a two-wire bus, as shown in Fig. 1.
@@ II-A CAN Protocol
Figure 1: CAN protocol frame with 8-byte payload. The full structure is presented for comprehensive understanding, with a specific focus on the ID and payload fields for our analysis.
Each CAN frame contains multiple fields; however, in the context of this work, we focus on two: the 11-bit arbitration ID and the up-to-8-byte data field. The ID determines message priority, with lower values corresponding to higher priority. The data field holds signal-level information such as speed or engine temperature. Messages with the same ID are transmitted periodically with a fixed payload structure, enabling the modeling of temporal and structural relationships among signals. For example, a powertrain ECU may broadcast a frame with ID 0x102 at 20 Hz containing real-time values for speed and odometer. These periodic patterns are typically defined using a CAN database (DBC) file, which decodes binary payloads into meaningful physical signals. This representation is particularly useful for downstream ML-based IDS that leverage time series analysis and graph-based representations [Zhang2023].
Beyond the protocol’s operational details, it is crucial to understand the security landscape and the various types of attacks targeting the CAN bus. Our approach targets multiple tiers of CAN bus threats, which are typically classified into three categories based on adversarial intent [verma2022addressing]:
  • Fabrication attacks: Malicious messages are injected onto the bus by a rogue ECU, without disrupting normal traffic from other ECUs [gazdag2023crysys].
  • Suspension attacks: A legitimate ECU is disabled or disconnected, causing its expected messages to vanish temporarily [rajapaksha2024canmirgu].
  • Masquerade attacks: The most stealthy and severe—these suppress a legitimate ECU and impersonate it by replaying spoofed frames at realistic timing and content patterns [Moriano:2022:Signal:Based:IDS].
Note that while this study focuses on masquerade attacks due to their stealth and impact, the proposed framework is also applicable to fabrication and suspension attacks, as these introduce structural and temporal deviations captured by MSGs.
II-B CAN Message Sequence Graph (MSG)
@@
The MSG is a crucial tool for analyzing the behavior of the CAN [Jedh2021]. It captures the sequence pattern of messages, reflecting their normal sequence of messages in CAN. Nodes in this graph correspond to unique CAN identifiers (IDs), denoting different ECUs or functions within the vehicle network. Edges map the sequence of messages between these IDs illustrating how information flows within the system. The MSG models regular patterns of communication between ECUs within specified sliding windows, allowing for detailed observation of both normal and potentially malicious activities. Typical sequence patterns found during various vehicle operations, such as parking or acceleration, can be captured by the MSG. These patterns form the basis for detecting deviations that may indicate cybersecurity threats like fabrication and suspension attacks.
Constructing an MSG involves parsing time-stamped CAN messages into a graph structure that captures the sequence and frequency of CAN ID transitions. This highlights communication patterns typical of normal vehicle operations—such as fuel delivery and speed coordination during acceleration—and enables detection of disruptions indicative of attacks. As shown in Section IV-A, MSGs visualize these dynamics, supporting robust analysis of ECU interactions and vehicular communication integrity.